# Robust inference of gravitational wave source parameters in the presence of noise transients using normalizing flows

### v1

@@ II.3 Analytical approach
The Kolmogorov-Smirnov (KS) test Lopes (2011) is a statistical method that can be used to test the goodness of fit between a sample distribution and a reference distribution. In the context of GW parameter estimation, the KS test can be employed to assess the agreement between the estimated parameter distribution and a theoretical or expected distribution Sidery et al. (2014); Veitch et al. (2015); Biwer et al. (2019); Thrane and Talbot (2020). The KS test first calculates the cumulative distribution function (CDF) of the two distributions. Then it evaluates the distance $D_{\rm{n}}$ between the two distributions based on the CDF, which can be written as
@@
where $F_{n}(x)$ is a specified continuous distribution function from a random sample that a set of random variables $X_{i}$, $i=1,...,n$. $F_{n}(x)$ and $F(x)$ respectively represent the CDFs of the two distributions under validation. The resulting $D_{n}$ is also known as the KS statistic. The p-value, calculated based on the KS statistic, is commonly employed as a straightforward measure to assess the similarity or dissimilarity between two distributions. The p-value is commonly compared to the significance level $\alpha$, and if the p-value is less than $\alpha$, it indicates that the two distributions are considered statistically different. Here, the significance level $\alpha$ has taken a value of 0.05. If the p-value obtained from the KS test is less than 0.05, it indicates that the two distributions are statistically different, and the hypothesis of their similarity can be rejected.
In this work, we compare two distributions: the ideal distribution and the one-dimensional distribution generated by the normalizing flow network. Since the percentile should be uniformly distributed between 0 and 1, the ideal distribution is uniform. The other distribution consists of the probabilities corresponding to the confidence intervals in which the true value of each sample lies among the simulated 200 samples. Our initial hypothesis assumes that the two distributions are the same distribution, which means the posterior inference of the parameters given by the normalizing flow network is reasonable. Then the p-value is compared with the significance level $\alpha$ and if the p-value was less than $\alpha$, we considered the original hypothesis to be invalid.
The Jensen–Shannon divergence (JSD) Lin (1991) is used to compare the difference between two probability distributions and measure the similarity of two distributions. It is achieved by taking a weighted average of the probability density functions of the two distributions, given by
@@ III Results and discussion
Before proceeding with further analysis, we conducted tests using normalizing flows to assess the rationality of parameter inference in the presence of glitches. We injected glitches into the simulated GW data and performed parameter inference using the normalizing flow and bilby dynesty respectively. The results are shown in Fig. 3. It can be seen that the parameter posterior distribution obtained by the normalizing flow is closer to the injection values of parameters than the results inferred by the bilby dynesty. This shows that the normalizing flow is superior to the conventional bilby dynesty for processing GW data contaminated by glitches, which is the basis of our subsequent work.
@@
We simulate 200 events to validate the robustness of the normalizing flow network’s inference for arbitrarily selected parameters. By injecting 200 simulated GW waveforms, along with the added glitches, we measured the frequency of the true parameter over a range of confidence levels. We calculate the cumulative probability of each parameter posterior’s percentile at the injection values of parameters, then plot the CDF for the joint distribution of the source parameters in Fig. 4. Ideally, the injected CDF of the parameters should exhibit a diagonal pattern, as the CDF represents the cumulative sum of individual event distributions. The proximity of the curve to the diagonal indicates the accuracy of parameter inference. A curve that closely aligns with the diagonal signifies superior performance in accurately estimating the parameters. The legend displays p-values obtained from the posterior distributions for each method.
@@
Based on our assumption that p-values less than 0.05 are deemed unreasonable for the model to support true posterior inference, the p-values provide a quantitative measure of the agreement between the normalizing flow model cumulative distribution and the expected theoretical distribution. The normalizing flow employed in our study demonstrates superior performance compared to the bilby dynesty when it comes to accurately learning the true posterior distribution, particularly in scenarios involving glitch-contaminated GW data.
@@
Figure 3: Parameter posterior plot of injected signal in Gaussian noise with glitch SNR of 8. Here the blue and orange lines show the inference of the joint posterior distribution of the parameters by the normalizing flow and bilby dynesty, respectively. One-dimensional histograms of the posterior distribution of each parameter in both methods are plotted along the diagonal. The injection values of the simulated signal are indicated by the black vertical and horizontal lines.
@@
Figure 4: Cumulative distribution of the quantile of which the true value lies for marginalized distribution. The curves were constructed using the 200 samples and the dashed black diagonal line indicates the ideal result. Shaded areas indicate 95% confidence intervals from the ideal distribution with the same number of events as the injection samples. The legend shows p-values obtained by the normalizing flow and bilby dynesty.
@@
Figure 5: Variation of the difference in the mean and standard deviation of inferred GW source parameter $d_{\rm L}$ with increasing glitch SNR. The dark and light gray lines represent the results from five independent experiments for the normalizing flow and the bilby dynesty, respectively. The arrows represent the direction of increasing glitch SNR. The blue and orange lines represent the average results from five independent experiments. The dark and light shades of the blue and orange lines respectively represent glitch SNR less than 8 and greater than 8.
@@
We selected the $d_{\rm L}$ parameter for demonstration purposes. As the glitch SNR increases, the mean and standard deviation obtained by the normalizing flow tends to be centered around the injected values when the glitch SNR is relatively low. However, they gradually deviate from the injected values as the glitch SNR increases. In contrast, the bilby dynesty continues to deviate from the injected values persistently as the glitch SNR increases. This also demonstrates the results above, specifically in Fig. 3, where the normalizing flow provides a wider range of posterior distributions as the injected values increase, while the bilby dynesty shows more deviation from the injected values. Indeed, this also highlights the robustness of the normalizing flow in handling glitches. The normalizing flow exhibits a certain level of stability and resilience in the presence of glitches, as evidenced by its ability to provide posterior distributions that are closer to the true values compared to the bilby dynesty.
@@ III.1 Different numbers of detectors
Figure 6: P-values of different glitch intensities with different number of detectors. The legend provides information about the colors and labels associated with the curves representing different detectors. Specifically, “LH” represents the LIGO network, “LHV” represents the LIGO and Virgo network, and “LHVK” represents the LIGO, Virgo, and KAGRA network Abbott et al. (2018b).
Figure 6 illustrates the p-values obtained parameter inference when glitches were injected into the GW data from different numbers of detectors, respectively. It depicts the distribution of p-values error bars for 20 sets of 200 GW data with varying detectors as the injected glitch SNR increases. These error bars are derived from the $68.3\%$ confidence intervals of the values provided by each group. The curves of the detectors exhibit trend an upward and then a downward trend. The normalizing flow can exhibit overconfident posterior approximations, leading to underestimated parameter variances Hermans et al. (2022). However, as shown in Fig. 5, it can be observed that when the glitch SNR is low, the standard deviation of the parameters tends to be larger. This behavior partially offsets the excessive confidence in the normalizing flow, leading to an increase in p-values in parameter inference calculations when glitches are absent. As a result, the plot exhibits a trend of initially increasing and then decreasing.
@@ III.3 Non-stationary noise
The parameters of glitch that we explore are those used in the simulation of glitch: frequency $f$, intensity $h$ and signal duration $Q$ Canton et al. (2014). Figure 9 shows the p-values plot of the normalizing flow network affected by the glitch parameter, drawn according to the KS test. We vary the glitch’s frequency, strength, and duration and ensure they are within a reasonable range. The horizontal and vertical coordinates on the graph are the two combinations of the glitch’s parameters. According to our assumption that the p-value is less than 0.05, the inference of network parameters is unreasonable. It can be seen in Fig. 9 that each parameter of glitch has an impact on the accuracy of the network’s inference of GW parameters. Specifically, as the duration of the glitch increases, along with a decrease in frequency and an increase in amplitude, the adverse effect on the rationality of parameter inference becomes more significant. Taking into account the upper limit of the GW signals frequency range, which is 200 to 400 Hz, and the lower limit of 20 Hz, it is evident that the impact of glitches on the inference of parameter inference becomes increasingly significant as more glitches overlap in both time and frequency with the GW signals.
@@ IV Conclusion
When GW data is contaminated by unknown non-Gaussian noise, it is possible to directly perform parameter inference using a normalizing flow network trained only on Gaussian noise. For unknown and unmodeled glitches, if the glitch SNR is within an appropriate range, the robustness of the normalizing flow network ensures reasonable parameter inferences. Additionally, we investigate the reasons behind the differences in the posterior distribution of source parameters observed between the normalizing flow and Bayesian inference methods, providing detailed explanations.
@@ Acknowledgements.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. We thank Shang-Jie Jin and Wan-Ting Hou for the helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).

### v2

@@ II.3 Analytical approach
The Kolmogorov-Smirnov (KS) test Lopes (2011) is a statistical method that can be used to test the goodness of fit between a sample distribution and a reference distribution. In the context of GW parameter estimation, the KS test can be employed to assess the agreement between the estimated parameter distribution and a theoretical or expected distribution Sidery et al. (2014); Veitch et al. (2015); Biwer et al. (2019); Thrane and Talbot (2020). The KS test first calculates the cumulative distribution functions (CDFs) of the two distributions. Then it evaluates the distance $D_{\rm{n}}$ between the two distributions based on the CDFs, which can be written as
@@
where $F_{n}(x)$ is a specified continuous distribution function from a random sample that a set of random variables $X_{i}$, $i=1,...,n$. $F_{n}(x)$ and $F(x)$ respectively represent the CDFs of the two distributions under validation. The resulting $D_{n}$ is also known as the KS statistic. The p-value, calculated based on the KS statistic, is commonly employed as a straightforward measure to assess the similarity or dissimilarity between two distributions. The p-value is commonly compared to the significance level. If the p-value is less than the significance level, it indicates that the two distributions are considered statistically different. Here, the significance level has taken a value of 0.05. If the p-value obtained from the KS test is less than 0.05, it indicates that the two distributions are statistically different, and the hypothesis of their similarity can be rejected.
In this work, we compare two distributions: the ideal distribution and the one-dimensional distribution generated by the normalizing flow network. Since the percentile should be uniformly distributed between 0 and 1, the ideal distribution is uniform. The other distribution consists of the probabilities corresponding to the confidence intervals in which the true value of each sample lies among the simulated 200 samples. Our initial hypothesis assumes that the two distributions are the same distribution, which means the posterior inference of the parameters given by the normalizing flow network is reasonable. Then the p-value is compared with the significance level of 0.05; if the p-value is less than 0.05, we consider the original hypothesis invalid.
The Jensen–Shannon divergence (JSD) is used to compare the difference between two probability distributions and measure the similarity of two distributions Lin (1991); Rao (1987). It is achieved by taking a weighted average of the probability density functions of the two distributions, given by
@@ III Results and discussion
Before proceeding with further analysis, we conducted tests using the normalizing flow to assess the rationality of parameter inference in the presence of glitches. We injected glitches into the simulated GW data and performed parameter inference using the normalizing flow and bilby dynesty respectively. The results are shown in Fig. 3. It can be seen that the parameter posterior distribution obtained by the normalizing flow is closer to the injection values of parameters than the results inferred by the bilby dynesty. This shows that the normalizing flow is superior to the conventional bilby dynesty for processing GW data contaminated by glitches, which is the basis of our subsequent work.
@@
We simulate 200 events to validate the robustness of the normalizing flow network’s inference for arbitrarily selected parameters. By injecting 200 simulated GW waveforms, along with the added glitches, we measured the frequency of the true parameter over a range of confidence levels. We also simulated 200 events without glitches. We calculate the cumulative probability of each parameter posterior’s percentile at the injection values of parameters, then plot the CDF for the joint distribution of the source parameters in Fig. 4. Ideally, the injected CDF of the parameters should exhibit a diagonal pattern, as the CDF represents the cumulative sum of individual event distributions. The proximity of the curve to the diagonal indicates the accuracy of parameter inference. A curve that closely aligns with the diagonal signifies superior performance in accurately estimating the parameters. The legend displays p-values obtained from the posterior distributions for each method.
@@
Based on our assumption that p-values less than 0.05 are deemed unreasonable for the model to support true posterior inference, the p-value provide a quantitative measure of the agreement between the normalizing flow model cumulative distribution and the expected theoretical distribution. With only Gaussian noise, both the normalizing flow and bilby dynesty methods perform well. But the normalizing flow employed in our study demonstrates superior performance compared to the bilby dynesty when it comes to accurately learning the true posterior distribution, particularly in scenarios involving glitch-contaminated GW data.
@@
Figure 3: Parameter posterior corner plot of injected signal in Gaussian noise with glitch SNR of 8. Here the blue and orange lines show the inference of the joint posterior distribution of the parameters by the normalizing flow and bilby dynesty, respectively. One-dimensional histograms of the posterior distribution of each parameter in both methods are plotted along the diagonal. The injection values of the simulated signal are indicated by the black vertical and horizontal lines.
@@
Figure 4: Cumulative distribution of the quantile of which the true value lies for marginalized distribution. The curves were constructed using the 200 samples and the dashed black diagonal line indicates the ideal result. Shaded areas indicate 95% confidence intervals from the ideal distribution with the same number of events as the injection samples. The legend shows p-values obtained by the normalizing flow and bilby dynesty. The triangular line represents p-values without glitches.
@@
Figure 5: Variation of the difference in the mean and standard deviation of inferred GW source parameter $d_{\rm L}$ with increasing glitch SNR. The dark and light gray lines represent the results from five independent experiments for the normalizing flow and the bilby dynesty, respectively. The arrows represent the direction of increasing glitch SNR. The blue and orange lines represent the average results from five independent experiments. Circles and squares indicate glitch SNR below and above 8, respectively.
@@
We selected the $d_{\rm L}$ parameter for demonstration purposes. As the glitch SNR increases, the mean and standard deviation obtained by the normalizing flow tend to be centered around the injected values when the glitch SNR is relatively low. However, they gradually deviate from the injected values as the glitch SNR increases. In contrast, the bilby dynesty continues to deviate from the injected values persistently as the glitch SNR increases. This also demonstrates the results above, specifically in Fig. 3, where the normalizing flow provides a wider range of posterior distributions as the injected values increase, while the bilby dynesty shows more deviation from the injected values. Indeed, this also highlights the robustness of the normalizing flow in handling glitches. The normalizing flow exhibits a certain level of stability and resilience in the presence of glitches, as evidenced by its ability to provide posterior distributions that are closer to the true values compared to the bilby dynesty.
@@ III.1 Different numbers of detectors
Figure 6: P-values of different glitch intensities with different number of detectors. The legend provides information about the colors and labels associated with the curves representing different detectors. Specifically, “LH” represents the LIGO network, “LHV” represents the LIGO and Virgo network, and “LHVK” represents the LIGO, Virgo, and KAGRA network Abbott et al. (2018b). The black line represents a p-value of 0.05.
Figure 6 illustrates p-values obtained parameter inference when glitches were injected into the GW data from different numbers of detectors, respectively. It depicts the distribution of p-values error bars for 20 sets of 200 GW data with varying detectors as the injected glitch SNR increases. These error bars are derived from the $1\sigma$ confidence intervals of the values provided by each group. The curves of the detectors exhibit an upward trend and then a downward trend. The normalizing flow can exhibit overconfident posterior approximations, leading to underestimated parameter variances Hermans et al. (2022). However, as shown in Fig. 5, it can be observed that when the glitch SNR is low, the mean value of the parameters does not change significantly, while the standard deviation increases noticeably. This behavior partially offsets the excessive confidence in the normalizing flow, leading to an increase in p-values in parameter inference calculations when glitches are absent. Therefore, when the glitch is small, the p-value will tend to increase. Since the p-values are bounded at 1, the expansion of the upper bounds of the error bars is constrained, while a more pronounced rise is observed in the lower bounds. This interplay between the inherent overconfidence of the network and the additional variance introduced by glitches leads to the tightening effect in Fig. 6 when the glitch SNR reaches 6.
@@ III.3 The SNR of the glitch
The parameters of glitch that we explore are those used in the simulation of glitch: $f_{0}$, $h_{0}$ and $Q$. Figure 9 shows the p-values plot of the normalizing flow network affected by the glitch parameter, drawn according to the KS test. We vary the glitch’s frequency, strength, and duration and ensure they are within a reasonable range. The horizontal and vertical coordinates on the graph are the two combinations of the glitch’s parameters. According to our assumption that the p-value is less than 0.05, the inference of network parameters is unreasonable. It can be seen in Fig. 9 that each parameter of glitch has an impact on the accuracy of the network’s inference of GW parameters. Specifically, as the duration of the glitch increases, along with a decrease in frequency and an increase in amplitude, the adverse effect on the rationality of parameter inference becomes more significant. Taking into account the upper limit of the GW signals frequency range, which is 200 to 400 Hz, and the lower limit of 20 Hz, it is evident that the impact of glitches on the inference of parameter inference becomes increasingly significant as more glitches overlap in both time and frequency with the GW signals.
@@ IV Conclusion
When GW data are contaminated by unknown non-Gaussian noise, it is possible to directly perform parameter inference using a normalizing flow network trained only on Gaussian noise. For unknown and unmodeled glitches, if the glitch SNR is within an appropriate range, the robustness of the normalizing flow network ensures reasonable parameter inferences. Additionally, we investigate the reasons behind the differences in the posterior distribution of source parameters observed between the normalizing flow and Bayesian inference methods, providing detailed explanations.
@@ Acknowledgements.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. We thank Shang-Jie Jin and Wan-Ting Hou for the helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).
Figure 10: The effect of $t_{0}$ on the inference of network parameters. The variation of p-values over different $t_{0}$. The blue error bar represents the $1\sigma$ confidence intervals for 20 sets of results. The black line represents a p-value of 0.05.
Figure 11: The effect of $\Delta t$ on the inference of network parameters. The p-values of parameter inference under different $\Delta t$. The blue curves illustrate the $1\sigma$ confidence bounds. The black line represents a p-value of 0.05.
Appendix A Glitch time parameter
In this Appendix, we examine the impact of glitch central time $t_{0}$ and its interval with the GW signal $\Delta t$ on the inference of network parameters.
To mitigate the impact of $\Delta t$, $t_{0}$ is uniformly distributed in (0.7,0.9) s. Simultaneously, the GW signal is uniformly varied within a range of 0.1 s around, with adjustments ensuring its time remains between 0.7 and 0.9 s. The results are presented in Fig. 10. Figure 10 illustrates the variation in p-values for parameter inference using the normalizing flow network as $t_{0}$ changes. It depicts error bar for the $1\sigma$ confidence intervals for 20 sets of results. It shows that with the change in $t_{0}$, the variation of p-values exhibits a relatively smooth trend overall, without significant upward or downward trends. This suggests that a single change in $t_{0}$ does not significantly affect parameter inference robustness.
To study $\Delta t$ between the glitch and GW signal, we varied $\Delta t$ in $(-0.1,0.1)$ s. Glitch times were generated from the GW signal time and $\Delta t$ to assess its impact on parameter robustness inference. The results are shown in Fig. 11. Figure 11 displays the p-value for network parameter inference as $\Delta t$ varies. The curves in Fig. 11 also represent error bar curves for $20$ sets of data. It can be observed that when $\Delta t$ is relatively large, indicating glitches occurring either before or after the signal, parameter inference remains within a reasonable range. However, when $\Delta t$ is 0 s, indicating overlapping glitches and GW signals, the network parameter inference results are poorer, which is consistent with Ref. Macas et al. (2022). This indicates that $\Delta t$ influences network parameter inference, with a significant impact when they overlap.
Since a single change in $t_{0}$ does not affect network parameter inference and the overlap between glitches and GW signals has a significant impact, we fixed the glitch time at 0.8 s in this work. This ensures that $\Delta t$ is uniformly distributed in $(-0.1,0.1)$ s, mitigating its influence on subsequent discussions.